# Myopically Verifiable Probabilistic Certificates for Safe Control and Learning

### v1

@@ I Introduction
Autonomous control systems must make safe control decisions in real-time and in the presence of uncertainties. Many techniques have been developed for deterministic systems with bounded uncertainties that can assure long-term safety with efficient computation [1, 2, 3]. These works often leverage set-invariance-based approaches [4, 5, 6] to transform long-term design specifications into sufficient myopic conditions. However, when a system has uncertainty with unbounded support (e.g., Gaussian noise), imposing set invariance on the state space for short-term future can no longer guarantee long-term safety. This is because even if the set invariance is satisfied with high probability at each time, uncertainty and risk may still accumulate over time.
@@
Consequently, there are stringent trade-offs between ensuring safety for longer time horizons and reducing computation burden. For example, robust and stochastic barrier functions can be efficiently evaluated [7, 8, 9], but are insufficient to ensure long-term guarantees. On the other hand, conditions to control the probability of continuously satisfying set invariance conditions for a long period of time cannot be converted into myopic conditions. Similarly, reachability-based approaches (e.g., probabilistic reachability) often come with heavy computation to propagate uncertainties over time [10, 11, 12, 13, 14, 15]. These drawbacks pose significant challenges to control and learning methods for stochastic systems operating under latency-critical environments.
@@
To address such challenges, in this paper, we propose a novel technique, termed probabilistic invariance. This technique is inspired by set invariance, but it defines invariance on the probability space, rather than on the state space. In particular, the probabilistic invariance technique allows one to efficiently evaluate probability about scenarios of long-term trajectories, such as probabilistic reachability. As such, this technique inherits the computational efficiency of the invariance-based approaches and the long-term guarantees of the reachability-based approaches. An infinitesimal future trajectory of a stochastic dynamical system provides little information about its future invariant sets on the state space. In comparison, our key insight is that its probabilistic analog, namely an infinitesimal future value of a long-term probability, does provide explicit computable information about invariant sets in the space of probability values. Building upon this insight, we characterize myopic conditions that can ensure conditions on long-term probability at all times.
@@
Leveraging this feature, we first characterize myopic conditions for controlling two types of long-term probabilities to stay within a desirable range. The first type is the probability of forward invariance, which can be used to represent the probabilities of exiting certain sets (e.g., safe regions) within a time interval. The second type is the probability of forward convergence, which can be used to represent the probabilities of reaching or recovering to certain sets. For affine control systems, these conditions can be converted into linear control constraints, and thus can be easily integrated into convex/quadratic programs. Then, we apply the probabilistic invariance technique to safe control and learning problems. In the context of control, the proposed technique equips the nominal controllers, which are allowed to be neural networks and black-box, with long-term safety. In the context of learning, the proposed technique minimally changes the training procedures while ensuring the long-term safety of the control policies during and after training.
II Related Work
II-A Safety Certificate for Stochastic Systems
@@
For deterministic systems with bounded uncertainties, previous work has developed various techniques for safe control (see review articles such as [3, 1, 16]), and the combination of set invariance and reachability-based techniques has also been studied in [17]. As this paper considers stochastic systems, our review below focuses on existing approaches for stochastic systems. We approximately classify these approaches into three main types based on their choice of tradeoffs: long-term safety with heavy computation; myopic safety with low computation; and long-term conservative safety with low computation.
@@
Extensive literature exists that considers the distributions of the system state in the long term. Reachability-based techniques characterize the probability that states reach or avoid certain regions over a long period of time [10, 11, 12, 13, 14, 15], which are useful to evaluate the long-term safety of actions. However, these techniques often come with significant computational costs. The cause is two-fold. First, possible trajectories often scale exponentially with the length of the outlook time horizon. Second, rare events are more costly to sample and estimate than nominal events. Compared to these techniques, the proposed method only myopically evaluates the immediate future, and the safety constraints can be given in closed forms for affine control systems. These features help reduce real-time computation.
@@
On the other hand, extensive work has been done to develop efficient controllers in latency-critical stochastic systems. Given the hard tradeoffs between time horizon vs. computation, many of these techniques impose set invariance on short-term or infinitesimal future states. For example, stochastic control barrier functions use a sufficient condition to ensure that the state moves within the tangent cone of the safe set on average [18]. The probabilistic barrier certificate ensures that conditions based on control barrier functions are met with high probability [8, 9]. The myopic nature of these methods achieves a significant reduction in computational cost. However, they can result in unsafe behaviors on a longer time horizon due to the accumulation of tail probabilities of hazardous events. In contrast, the proposed technique provides long-term safe probability guarantees by finding probabilistic invariance conditions that directly control the likelihood of accumulating tail events.
@@
II-B Integration with Learning-based Techniques
@@
The safety certificates mentioned above (e.g., [3, 18, 8, 9]) are commonly used to certify or modify nominal controllers. The nominal controllers are often allowed to be either black-box models or represented by neural networks. As stated above, in stochastic systems, when safety certificates are formulated using myopic conditions, they do not necessarily ensure long-term safety. When they are constructed based on approximations or designed to be safe in worst-case uncertainties, the system performance may degrade significantly as the size of uncertainties increases. On the other hand, the proposed certificate can be used similarly to existing safety certificates but ensures long-term safety using myopic controllers and exhibits graceful degradation for increasing uncertainties. [^1] For example, in its application to extreme driving, the former can be observed from [24, Fig. 2], and the latter from [24, Fig. 6]. Alternatively, safety considerations can be captured in rewards and constraints in model predictive control (MPC) [25, 26]. When safety specifications for future states are directly expressed in the reward and constraint functions, the MPC outlook time horizon needs to cover such future states. In comparison, the proposed technique can be used to construct MPC constraints in a way that ensures safety for a longer period than the MPC outlook horizon.
@@
There exists extensive literature on safe reinforcement learning, as reviewed in [27, 28]. Some approaches account for safety specifications in rewards or constraints [29, 30, 31, 32, 33] or design/learn certain functions related to safe conditions [34, 35, 35, 36, 37, 38]. Examples of these functions are control barrier functions [34], safety index [35], barrier certificate [36], rewards and constraints [37, 38]. These methods aim to learn optimal and safe policies but may not guarantee safety in the initial learning phase. Other approaches impose additional structures on control policies [39, 40, 41]. For example, safety filters (e.g., Lyapunov conditions, control barrier functions, typically defined by prior knowledge of system models) are attached to learning-based controllers, and can often ensure safety or stability for deterministic systems or bounded uncertainties. Similarly, the proposed technique can be used to construct a safety filter for learning-based controllers. When used as a safety filter, it can control, both during and after training, the probability of forward invariance for a given duration or forward convergence within a given horizon for stochastic nonlinear systems.
III Preliminary
@@
Let $\mathbb{R}$, $\mathbb{R}_{+}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{m\times n}$ be the set of real numbers, the set of non-negative real numbers, the set of $n$-dimensional real vectors, and the set of $m\times n$ real matrices, respectively. Let $x[k]$ be the $k$-th element of vector $x$. Let $X_{a:b}$ denote the set $\{X_{i}:i\in\{a,a+1,\cdots,b\}\}$, where $a,b\in\mathbb{Z}$ and $a\leq b$. Let $f:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ represent that $f$ is a mapping from space ${\mathcal{X}}$ to space ${\mathcal{Y}}$. Let $\mathbb{1}({\mathcal{E}})$ be an indicator function, which takes $1$ when condition ${\mathcal{E}}$ holds and 00 otherwise. Let $\mathbf{0}_{m\times n}$ be an $m\times n$ matrix with all entries 00. Given events ${\mathcal{E}}$ and ${\mathcal{E}}_{c}$, let $\mathbb{P}({\mathcal{E}})$ be the probability of ${\mathcal{E}}$ and $\mathbb{P}({\mathcal{E}}|{\mathcal{E}}_{c})$ be the conditional probability of ${\mathcal{E}}$ given the occurrence of ${\mathcal{E}}_{c}$. Given random variables $X$ and $Y$, let $\mathbb{E}[X]$ be the expectation of $X$ and $\mathbb{E}[X|Y=y]$ be the conditional expectation of $X$ given $Y=y$. We use upper-case letters (e.g., $Y$) to denote random variables and lower-case letters (e.g., $y$) to denote their specific realizations.
@@
IV Problem Statement
Here, we introduce the control system in subsection IV-A, define the safety specification in subsection IV-B, and state the controller design goals in subsection IV-C.
IV-A Control System Description
@@
where $X_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the system state, $U_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the control input, and $W_{t}\in\mathbb{R}^{\xi}$ captures the system uncertainties. Here, $X_{t}$ can include both the controllable states of the system and the uncontrollable environmental variables such as moving obstacles. We assume that $W_{t}$ is the standard Wiener process with 00 initial value, i.e., $W_{0}=0$. The value of $\sigma(X_{t})$ is determined based on the size of uncertainty in unmodeled dynamics, environmental variables and noise.
The control action $U_{t}$ is determined at each time by the control policy. We assume that accurate information of the system state can be used by the control policy. The control policy is composed of a nominal controller and additional modification scheme to ensure the safety specifications illustrated in subsection IV-B. The nominal controller is represented by
@@
where function $N$ lies in a class of stochastic or deterministic functions mapping $\mathcal{X}$ to $\mathcal{U}$, i.e., $N\in\mathcal{N}_{s}\cup\mathcal{N}_{d}$. Here, $\mathcal{N}_{s}$ denotes the class of stochastic functions, $\mathcal{N}_{d}$ denotes the class of stochastic functions, and we use $\mathcal{N}$ to denote the function class when it can be either. The design of $N$ does not necessarily account for the safety specifications defined below and can be represented by neural networks. To adhere to the safety specifications, the output of the nominal controller is then modified by another scheme. The overall control policy involving the nominal controller and the modification scheme is represented by
@@
where $N^{\prime}:\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}^{m}$ is a mapping from the current state $X_{t}$, safety margin $L_{t}$, and time horizon $T_{t}$ to the current control action $U_{t}$. The formal definition of $L_{t}$ and $T_{t}$ will be given later in the section. The policy of the form (4) assumes that the decision rule is time-invariant. [^2] The functions $N$ and $N^{\prime}$ do not change over time. This policy is also assumed to be memory-less in the sense that it does not use the past history of the state $\{X_{\tau}\}_{\tau<t}$ to produce the control action $U_{t}$. The assumption for memory-less controller is reasonable because the state evolution $dX_{t}$ of system (2) only depends on the current system state $X_{t}$ as $f(X_{t})$, $g(X_{t})$, and $\sigma(X_{t})$ are time-invariant functions of the system state. We restrict ourselves to the settings when $f$, $g$, $\sigma$, $N$, and $N^{\prime}$ have sufficient regularity conditions such that the closed loop system of (2) and (4) has unique strong solutions. Conditions required to have a unique strong solution can be found in [42, Chapter 1], [43, Chapter II.7] and references therein.
@@
to denote the $L$-super level set of $\phi$ (the set with safety margin $L$). Accordingly, we use $\text{int}\,{{\mathcal{C}}(L)}=\{x\in\mathbb{R}^{n}:\phi(x)>L\}$ to denote the interior of the safe set, ${\mathcal{C}}(L)^{c}=\{x\in\mathbb{R}^{n}:\phi(x)<L\}$ to denote the unsafe set, $\partial{\mathcal{C}}(L)=\{x\in\mathbb{R}^{n}:\phi(x)=L\}$ to denote the boundary of $L$ super level set.
IV-B Safety Specifications
@@ IV-B1 Forward Invariance
for some time interval $[t,t+T_{t}]$ conditioned on an initial condition $x\in{\mathcal{C}}(L)$. Probability (7) can be computed from the distribution of the following two random variables:333 These random variables are previously introduced and analyzed in [44].
$\Phi_{x}(T_{t})$ $:=\inf\{\phi(X_{\tau})\in\mathbb{R}:\tau\in[0,T_{t}],X_{0}=x\},$ (8)
$\Gamma_{x}(L_{t})$ $:=\inf\{\tau\in\mathbb{R}_{+}:\phi(X_{\tau})<L_{t},X_{0}=x\}.$ (9)
Here, $\Phi_{x}(T_{t})$ is the worst-case safety margin from the boundary of the safe set $\partial{\mathcal{C}}(0)$ during $[0,T_{t}]$, and $\Gamma_{x}(L_{t})$ is the time when the system exit from ${\mathcal{C}}(L_{t})$ for the first time. We can rewrite (7) using the two random variables (8) and (9) as
$=\>\mathbb{P}\left(X_{\tau}\in{\mathcal{C}}(L_{t}),\forall\tau\in[0,T_{t}]\ |\ X_{0}=x\right)$ (11)
$=\>\mathbb{P}(\Phi_{x}(T_{t})\geq L_{t})$ (12)
[…]
Here, equality (11) holds due to the time-invariant nature of the system and control policies.
@@
$P_{1}(x,L,T,\phi):=\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L),\forall\tau\in[t,t+T]\mid X_{t}=x\},$ (14)
@@
$P_{2}(x,L,T,\phi):=\sup_{N\in\mathcal{N}}\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L),$ (15)

### v2

@@ 1 Introduction
Autonomous control systems must make safe control decisions in real-time and in the presence of uncertainties. For example, autonomous driving cars need to produce specific pedal, brake and steering commands to ensure their safety on road while interacting with other road users. Many techniques have been developed for deterministic systems with bounded uncertainties that can assure long-term safety with efficient computation [1, 2, 3]. These works often leverage set-invariance-based approaches [4, 5, 6] to transform long-term design specifications into sufficient myopic conditions. However, when a system has uncertainty with unbounded support (e.g., Gaussian noise), imposing set invariance on the state space for short-term future can no longer guarantee long-term safety. This is because even if the set invariance is satisfied with high probability at each time, uncertainty and risk may still accumulate over time.
@@
Consequently, there are stringent trade-offs between ensuring safety for longer time horizons and reducing computation burden. For example, robust and stochastic barrier functions can be efficiently evaluated [7, 8], but are insufficient to ensure long-term guarantees [9]. On the other hand, conditions to control the probability of continuously satisfying set invariance conditions for a long period of time cannot be converted into myopic conditions [10]. Similarly, reachability-based approaches (e.g., probabilistic reachability) often come with heavy computation to propagate uncertainties over time [11, 12, 13, 14, 15, 16]. These drawbacks pose significant challenges to control and learning methods for stochastic systems operating under latency-critical environments.
@@
To address such challenges, in this paper, we propose a novel technique, termed probabilistic invariance. This technique is inspired by set invariance, but it defines invariance on the probability space, rather than on the state space. This technique inherits the computational efficiency of the invariance-based approaches and the long-term guarantees of the reachability-based approaches. An infinitesimal future trajectory of a stochastic dynamical system provides little information about its future invariant sets on the state space. In comparison, our key insight is that its probabilistic analog, namely an infinitesimal future value of a long-term probability, does provide explicit computable information about invariant sets in the space of probability values.
@@
Building upon this insight, we first characterize myopic conditions that can ensure conditions on long-term probability at all times. Specifically, the conditions can be used to constrain two types of long-term probabilities to stay within a desirable range. The first type is the probability of forward invariance, which can be used to represent the probabilities of exiting certain sets (e.g., safe regions) within a time interval. The second type is the probability of forward convergence, which can be used to represent the probabilities of reaching or recovering to certain sets. For affine control systems, these conditions can be converted into linear control constraints, and thus can be easily integrated into convex/quadratic programs. Then, we apply the probabilistic invariance technique to safe control and learning problems. In the context of control, the proposed technique equips the nominal controllers, which are allowed to be neural networks and black-box controllers, with long-term safety. In the context of learning, the proposed technique minimally changes the training procedures while ensuring the long-term safety of the control policies during and after training. The contributions of this paper are summarized below:
  • We propose a novel technique termed probabilistic invariance, which can be used to find myopic conditions with assured long-term guarantees (Theorem 1).
  • Building upon these conditions, we present probabilistic safety certificate to be used in real-time control (Section 5) and test its performance in simulated environments (Section 7).
  • We show how to integrate the probabilistic certificate with policy gradient and Q-learning (Section 6) and demonstrate their efficacy in simulation (Section 8).
The rest of the paper is organized as follows. We present related work in Section 2 and preliminaries in Section 3. We then present the problem formulation in Section 4. We introduce the proposed probabilistic safety certificate in Section 5, where we establish theoretical guarantees for long-term safety and provide practical computation methods for implementation. Afterwards, we present how the proposed safety certificate can be integrated with reinforcement learning in Section 6. We present experiment results in Sections 7 and 8 and finally we conclude the paper in Section 9.
2 Related Work
Safety Certificate for Stochastic Systems
@@ 2 Related Work
For deterministic systems with bounded uncertainties, previous work has developed various techniques for safe control (see review articles such as [3, 1, 17]), and the combination of set invariance and reachability-based techniques has been studied in [18]. As this paper considers stochastic systems, our review below focuses on existing approaches for stochastic systems. We approximately classify these approaches into three main types based on their choice of tradeoffs: long-term safety with heavy computation; myopic safety with low computation; and long-term conservative safety with low computation.
@@
Extensive literature considers the distributions of the system state in the long term. Reachability-based techniques characterize the probability that states reach or avoid certain regions over a long period of time [11, 12, 13, 14, 15, 16, 19], which are useful to evaluate the long-term safety of actions. However, these techniques often come with significant computational costs. The cause is two-fold. First, possible trajectories often scale exponentially with the length of the outlook time horizon. Second, rare events are more costly to sample and estimate than nominal events. Compared to these techniques, the proposed method only myopically evaluates the immediate future, and the safety constraints can be given in closed forms for affine control systems. These features help reduce real-time computation.
@@
On the other hand, extensive work has been done to develop efficient controllers in latency-critical stochastic systems. Given the hard tradeoffs between time horizon vs. computation, many of these techniques impose set invariance on short-term or infinitesimal future states. For example, stochastic control barrier functions use a sufficient condition to ensure that the state moves within the tangent cone of the safe set on average [20, 21, 22, 23]. The probabilistic barrier certificate ensures that conditions based on control barrier functions are met with high probability [10, 8, 24, 25]. The myopic nature of these methods achieves a significant reduction in computational cost. However, they can result in unsafe behaviors on a longer time horizon due to the accumulation of tail probabilities of hazardous events. In contrast, the proposed technique provides long-term safe probability guarantees by finding probabilistic invariance conditions that directly control the likelihood of accumulating tail events.
@@
Integration with Learning-based Techniques
@@
The safety certificates mentioned above (e.g., [3, 20, 10, 8]) are commonly used to certify or modify nominal controllers. The nominal controllers are often allowed to be either black-box models or represented by neural networks [30]. As stated above, in stochastic systems, when safety certificates are formulated using myopic conditions, they do not necessarily ensure long-term safety. When they are constructed based on approximations or designed to be safe in worst-case uncertainties, the system performance may degrade significantly as the size of uncertainties increases. On the other hand, the proposed certificate can be used similarly to existing safety certificates but ensures long-term safety using myopic controllers and exhibits graceful degradation for increasing uncertainties. [^1] For example, in its application to extreme driving, the former can be observed from [31, Fig. 2], and the latter from [31, Fig. 6]. Alternatively, safety considerations can be captured in rewards and constraints in model predictive control (MPC) [32, 33, 34]. When safety specifications for future states are directly expressed in the reward and constraint functions, the MPC outlook time horizon needs to cover such future states. In comparison, the proposed technique can be used to construct MPC constraints in a way that ensures safety for a longer period than the MPC outlook horizon.
@@
There exists extensive literature on safe reinforcement learning, as reviewed in [35, 36]. Some approaches account for safety specifications in rewards or constraints [37, 38, 39, 40, 41] or design/learn certain functions related to safe conditions [42, 43, 44, 45, 46]. Examples of these functions are control barrier functions [42], safety index [43], barrier certificate [44], rewards and constraints [45, 46]. These methods aim to learn optimal and safe policies but may not guarantee safety in the initial learning phase. Other approaches impose additional structures on control policies [47, 48, 49, 50, 51]. For example, safety filters (e.g., Lyapunov conditions, control barrier functions, typically defined by prior knowledge of system models) are attached to learning-based controllers, and can often ensure safety or stability for deterministic systems or bounded uncertainties. Similarly, the proposed technique can be used to construct a safety filter for learning-based controllers. When used as a safety filter, it can control, both during and after training, the probability of forward invariance for a given duration or forward convergence within a given horizon for stochastic nonlinear systems.
3 Preliminary
@@
Let $\mathbb{R}$, $\mathbb{R}_{+}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{m\times n}$ be the set of real numbers, the set of non-negative real numbers, the set of $n$-dimensional real vectors, and the set of $m\times n$ real matrices, respectively. Let $x[k]$ be the $k$-th element of vector $x$. Let $X_{a:b}$ denote the set $\{X_{i}:i\in\{a,a+1,\cdots,b\}\}$, where $a,b\in\mathbb{Z}$ and $a\leq b$. Let $f:{\mathcal{X}}\rightarrow{\mathcal{Y}}$ represent that $f$ is a mapping from space $\mathcal{X}$} to space $\mathcal{Y}$}. Let $\mathbb{1}({\mathcal{E}})$ be an indicator function, which takes $1$ when condition $\mathcal{E}$} holds and 0 otherwise. Let $\mathbf{0}_{m\times n}$ be an $m\times n$ matrix with all entries 0. Given events $\mathcal{E}$} and ${\mathcal{E}}_{c}$, let $\mathbb{P}({\mathcal{E}})$ be the probability of $\mathcal{E}$} and $\mathbb{P}({\mathcal{E}}|{\mathcal{E}}_{c})$ be the conditional probability of $\mathcal{E}$} given the occurrence of ${\mathcal{E}}_{c}$. Given random variables $X$ and $Y$, let $\mathbb{E}[X]$ be the expectation of $X$ and $\mathbb{E}[X|Y=y]$ be the conditional expectation of $X$ given $Y=y$. We use upper-case letters (e.g., $Y$) to denote random variables and lower-case letters (e.g., $y$) to denote their specific realizations.
@@
4 Problem Statement
Here, we introduce the control system in subsection 4.1, define the safety specification in subsection 4.2, and state the controller design goals in subsection 4.3.
4.1 Control System Description
@@
where $X_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ is the system state, $U_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ is the control input, and $W_{t}\in\mathbb{R}^{\xi}$ captures the system uncertainties. Here, $X_{t}$ can include both the controllable states of the system and the uncontrollable environmental variables such as moving obstacles. We assume that $W_{t}$ is the standard Wiener process with 0 initial value, i.e., $W_{0}=0$. The value of $\sigma(X_{t})$ is determined based on the size of uncertainty in unmodeled dynamics, environmental variables and noise.
The control action $U_{t}$ is determined at each time by the control policy. We assume that accurate information of the system state can be used by the control policy. The control policy is composed of a nominal controller and additional modification scheme to ensure the safety specifications illustrated in subsection 4.2. The nominal controller is represented by
@@
where function $N$ lies in a class of stochastic or deterministic functions mapping $\mathcal{X}$ to $\mathcal{U}$, i.e., $N\in\mathcal{N}_{s}\cup\mathcal{N}_{d}$. Here, $\mathcal{N}_{s}$ denotes the class of stochastic functions, $\mathcal{N}_{d}$ denotes the class of deterministic functions, and we use $\mathcal{N}$ to denote the function class when it can be either. The design of $N$ does not necessarily account for the safety specifications defined below and can be represented by neural networks. To adhere to the safety specifications, the output of the nominal controller is then modified by another scheme. The overall control policy involving the nominal controller and the modification scheme is represented by
@@
where $N^{\prime}:\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}^{m}$ is a mapping from the current state $X_{t}$, safety margin $L_{t}$, and time horizon $T_{t}$ to the current control action $U_{t}$. The formal definition of $L_{t}$ and $T_{t}$ will be given later in the section. The policy of the form (4) assumes that the decision rule is time-invariant. [^2] The functions $N$ and $N^{\prime}$ do not change over time. This policy is also assumed to be memory-less in the sense that it does not use the past history of the state $\{X_{\tau}\}_{\tau<t}$ to produce the control action $U_{t}$. The assumption for memory-less controller is reasonable because the state evolution $dX_{t}$ of system (2) only depends on the current system state $X_{t}$ as $f(X_{t})$, $g(X_{t})$, and $\sigma(X_{t})$ are time-invariant functions of the system state. We restrict ourselves to the settings where $f,g,\sigma,N,N$’ have sufficient regularity conditions such that the closed-loop system has a well-defined solution. For example, the system is well-defined when $f,g,\sigma,N,N$’ are bounded and globally Lipschitz [52], which are common assumptions usually satisfied in practice. More generally, conditions required to have a unique solution can be found in [53, Chapter 1], [54, Chapter II.7] and references therein.
@@
to denote the $L$-super level set of $\phi$ (the set with safety margin $L$). Accordingly, we use ${\mathcal{C}}(L)^{c}=\{x\in\mathbb{R}^{n}:\phi(x)<L\}$ to denote the unsafe set (complement of the safe set).
4.2 Safety Specifications
@@ 4.2.1 Forward Invariance
for some time interval $t,t+T_{t}]$ conditioned on an initial condition $x\in{\mathcal{C}}(L)$.
@@
$P_{1}(x,L,T,\phi)=\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L),\forall\tau\in[t,t+T]\mid X_{t}=x\},$ (8)
@@
$P_{2}(x,L,T,\phi)=\sup_{N\in\mathcal{N}}\mathbb{P}\{X_{\tau}\in{\mathcal{C}}(L),$ (9)